Compare the resistance of a 1.5-Amp interior light bulb of a car (operating off a 12-V battery) to the resistance of a 100-Watt bulb operating on a 110-volt household circuitry.
Answer #1

Given ,

Current of light bulb in car , I1 = 1.5 Amperes

Voltage of Bulb in car , V1 = 12 Volts

Using Ohm's Law , Resistance = Voltage / Current i.e., R = V/I

Resistance of 1.5 Ampere of 1.5 Amperes interior light bulb of car (operating at 12 Volts) , R1 = V1/I1 = 12/1.5 = 8 Ohms

Power of bulb in household , P2 = 100 Watts

Voltage operated in household , V2 = 110 Volts

We know that resistance = (Voltage)*(Voltage)/(Power) i.e., R = V^2/P

Resistance of house hold bulb , R2 = 110^2/100 = 12100/100 = 121 Ohms

Comparision of "Resistance of 1.5 Ampere of 1.5 Amperes interior light bulb of car (operating at 12 Volts)" to "resistance of a 100-Watt bulb operating on a 110-volt household circuitry." = R1/R2 = 8/121

Comparision of resistance values = R1/R2 = 8/121 = 0.06611570247

If you need more information please comment . Thank you